OpenStudy (anonymous):

i cant figure out the result of this integral using two methods: int ((a^2)+(x^2))^(-3/2)dx i used one method of substituting x with a*tan(theta). can anyone figure out another method? simpler is better. thanks

OpenStudy (dumbcow):

i cant think of it... doing the trig sub seems simple enough to me

OpenStudy (amoodarya):

use this substitution u=a * tan x then a^2+x^2=a^2(1+tan^2x )=a^2 sec^x

OpenStudy (dumbcow):

@amoodarya , read the whole post, that trig sub is already been mentioned

OpenStudy (anonymous):

its pretty easy Substitute x=\[a \tan \theta\] then dx=\[a \sec ^2\theta d \theta\] If the integral is I then now I= \[\int\limits_{}^{} \frac{ a \sec ^2\theta d \theta }{ (asec \theta)^3}\] I=\[\int\limits_{}^{}\frac{ 1 }{ a^2 }* \cos \theta d \theta\] I=\[\frac{ 1 }{ a^2 } \sin \theta + c\] I= \[\frac{ 1 }{ a^2} * \frac{ x }{ \sqrt{(a^2+x^2)} } + c\]

OpenStudy (anonymous):

Oh you asked for another method :P I didn't read the entire question. You can do this with integration by parts but trigonometric substitution is the shortest and simplest

OpenStudy (amoodarya):

you can use also u=sinh x so a^2+x^2=a^2(1+sinh^2x)=a^2 cosh^2x and du= cosh x dx

OpenStudy (anonymous):

thanks MayankD for the reply. How can I compute it by parts?

OpenStudy (anonymous):

Use \[\int\limits_{}^{}u*v*dx=u \int\limits_{}^{}v*dx - \int\limits_{}^{}[u' *\int\limits_{}^{}v*dx]dx\]

OpenStudy (anonymous):

how can I split the function to: \[\int\limits_{}^{}u*v*dx\] thankyou for the reply

OpenStudy (anonymous):

Take u as 1 and v as the rest of the function. Its a difficult integral to solve but with substitution you should be able to do it sooner or later.

OpenStudy (anonymous):

thankyou MayankD yes I agree that with substitution it is very simplier to solve the integral.